OpenStudy (anonymous):

A cyclist rides his bike at a speed of 33 feet per second. What is this speed in miles per hour? How many miles will the cyclist travel in 2 hours? In your computations, use the fact that 1 mile is equal to 5280feet. Do not round your answers

OpenStudy (anonymous):

\[\frac{ 33ft }{ s } * \frac{ 1 mi }{ 5280ft } * \frac{ 3600s }{ 1 hr. }* 2hrs.\] Notice that all units cancel out except for miles.

OpenStudy (anonymous):

45miles

OpenStudy (anonymous):

the speed is 45 right and the distance in 2 hrs is 90 ??

OpenStudy (anonymous):

The speed is 22.5mph. The distance is 45 miles

OpenStudy (anonymous):

88 ft/sec is 60 mph\[\frac{33}{88}60=\frac{45}{2}=22.5 \text{ mph} =45\text{ miles/2 hours}\]